How much NaCl (58.44 g/mol) would be needed to make 500 ml of a 2mM NaCl solution?
Answer #1

M = mol /V

V = 500 ml = 0.5 L

M = 2*10^-3 M

mol = MV = (2*10^-3)*0.5 = 0.001mol of NACl needed

mass = mol*MW = 0.001*58.44 = 0.05844 g of NaCl needed